Calculate the peak current in a 3.0?k? resistor connected to a 220-V rms ac source.


Ipeak= _________A
Answer #1

Vrms = V/sqrt(2)

So, V=Vrms*sqrt(2) = 220*sqrt(2) = 311.13V

Ipeak = V/R = 311.13/3000 = 0.1 A
Answer #2

Peak voltage = 220 * sqrt(3) V
So peak current = 220 * sqrt(3) V / 3 kOhm = 127.010 mA.